# Continuized Nesterov Acceleration for Non-Convex Optimization

### v1

@@ Abstract
In convex optimization, continuous-time counterparts have been a fruitful tool for analyzing momentum algorithms. Fewer such examples are available when the function to minimize is non-convex. In several cases, discrepancies arise between the existing discrete-time results — namely those obtained for momentum algorithms — and their continuous-time counterparts, with the latter typically yielding stronger guarantees. We argue that the continuized framework (even2021continuized), mixing continuous and discrete components, can tighten the gap between known continuous and discrete results. This framework relies on computations akin to standard Lyapunov analyses, from which are deduced convergence bounds for an algorithm that can be written as a Nesterov momentum algorithm with stochastic parameters. In this work, we extend the range of applicability of the continuized framework, e.g. by allowing it to handle non-smooth Lyapunov functions. We then strengthen its trajectory-wise guarantees for linear convergence rate, deriving finite time bounds with high probability and asymptotic almost sure bounds. We apply this framework to the non-convex class of strongly quasar convex functions. Adapting continuous-time results that have weaker discrete equivalents to the continuized method, we improve by a constant factor the known convergence rate, and relax the existing assumptions on the set of minimizers.
@@ 1.1 ODEs Momentum and their Limit for Algorithm Design
Note that $\{\tilde{x}_{k}\}_{k\in{\mathbb{N}}}$ reduces to gradient descent for the choice $\alpha_{k}=1$, $\forall k\in{\mathbb{N}}$. As previously mentioned, it is known to accelerate over gradient descent in many settings in convex optimization (suboydcandes; adr_qsc; aujdossrondPL; li2024linear). In particular, for $\mu$-strongly convex functions with $L$-Lipschitz gradient, $\{\tilde{x}_{k}\}_{k\in{\mathbb{N}}}$ generated by (4) properly parameterized ensures $f(\tilde{x}_{k})-f^{\ast}=\mathcal{O}((1-\sqrt{\mu/L})^{k})$ (nesterovbook). This should be compared with the bound $\mathcal{O}((1-\mu/L)^{k})$ using gradient descent, which can be significantly worse as typically $\mu\ll L$ (this is particularly true in the high dimensional setting). It is thus common to mention this gain of a squared dependence on $\mu/L$ as acceleration.
@@
Both rates correspond to linear convergence, with rates $\sqrt{\mu}$ and $\sqrt{\mu/L}$, respectively. In this sense, the continuous-time result is ”transferred” to the discrete setting. The $L$-Lipschitzness of the gradient is a key property to enable this transfer. In particular, the accelerated rate $(1-\sqrt{\mu/L})^{k}$ can be achieved if we take $s$ in (4) large enough, namely $s=1/L$. This has been extensively used in the convex setting where similar convergence results are obtained for $(x_{t})_{t\geq 0}$ and $\{x_{n}\}_{n\geq 0}$, see e.g. (suboydcandes; siegel2021accelerated; attouch2020firstorder; shi2022understanding; aujdossrondPL; hyppo; li2024linear). However, this approach may suffer from two drawbacks.
(i) Convergence properties do not transfer   The convergence properties of $(x_{t})_{t\geq 0}$ may not transfer to $\{x_{n}\}_{n\geq 0}$ when $f$ is not convex. In the case of $L$-smooth functions, if the total energy decreases when applied to the solution of the Heavy Ball equation, it is not the case for the discrete counterpart, see (jin2018accelerated, Section A.1). For functions with Lipschitz gradient and Hessian, a convergence result exists for the Heavy Ball equation, but needs additional restrictions to deduce a discrete counterpart, see the first version of (okamura2024primitive). In the case of strongly quasar convex functions, in order to select a stepsize large enough to achieve the best known convergence rate with $\{\tilde{x}_{k}\}_{k\in{\mathbb{N}}}$, the non-convexity has to be restricted (hermant2024study; guptanesterov). Such restrictions are not necessary when considering $(x_{t})_{t\geq 0}$. Using side by side comparisons of discrete and continuous Lyapunov functions, Section 4.2 in (hermant2024study) precises how non-convexity hurts the discretization process in this case.
@@ 1.2 A Promising Alternative: Continuized Nesterov
At $t=T_{i}$, the process ”jumps” by performing a gradient step
@@
$\left\{\begin{array}[]{ll}dx_{t}&=\eta_{t}(z_{t}-x_{t})dt-\gamma_{t}\nabla f(x_{t})dN(t)\\ dz_{t}&=\eta^{\prime}_{t}(x_{t}-z_{t})dt-\gamma^{\prime}_{t}\nabla f(x_{t})dN(t)\end{array}\right$. (14)
where $dN(t)=\sum_{k\geq 0}\delta_{T_{k}}(dt)$ is a Poisson point measure. Intuitively, since $\mathbb{E}\left[dN_{t}\right]=dt$, the mean process $(\mathbb{E}\left[x_{t}\right],\mathbb{E}\left[z_{t}\right])_{t\geq 0}$ is smooth, see Figure 1. In particular, if $\nabla f$ is linear, it can be written exactly in the form of (7). Consequently, individual realizations of the jump-driven process (14) provide a stochastic approximation of a much smoother expected trajectory. It implies that, for a given Lyapunov function $E_{t}:=\varphi(x_{t},z_{t},t)$, the analysis of $\mathbb{E}\left[E_{t}\right]$ is close to that arising in the standard continuous setting, such as in the study of (7).
@@
A non-convex application: strong quasar convexity   Strong quasar convexity is a non-convex relaxation of strong convexity, that has received increasing attention in recent years (hinder2020near; gower2021sgd; wang2023continuizedaccelerationquasarconvex; fu2023accelerated; alimisis2024characterization; hermant2024study; chenefficient2025; guptanesterov; lara2025delayed; de2025extending; farzin2025minimisation). Compared to gradient descent, better convergence bounds exist for Nesterov momentum when minimizing these functions. The first of these accelerated results mixes the momentum mechanism with a subroutine that solves an optimization sub-problem (quasarconvexold1; quasarconvexold2), reduced to a line-search procedure which adds a logarithmic factor in the complexity in (hinder2020near). Avoiding such a subroutine requires to restrict the class of function, excluding functions that are too non-convex (hermant2024study; guptanesterov). The continuized method has been applied to strongly quasar convex functions (wang2023continuizedaccelerationquasarconvex), recovering the convergence rate of the classical Nesterov momentum algorithm. Strikingly, it does so without requiring a line-search procedure or restricting the level of non-convexity. This indicates that the method is capable of achieving results that classical momentum algorithms have not been able to obtain so far, and we show in this work that it can go further.
@@ 1.3 Limits of the Framework and Contributions
(iii) Adapting results coming from the classic ODE framework literature, we improve the best known convergence rate by a constant factor for smooth strongly quasar convex functions. We further relax the existing assumption on the set of minimizers for this class of functions. Our results hold when considering unbiased stochastic estimates of the gradients, as long as a strong growth condition is verified.
@@ 2 Assumptions and Notations
We denote $\mathcal{X}^{\ast}:=\arg\min_{{\mathbb{R}}^{d}}f(x)$, $x^{\ast}$ any element of $\mathcal{X}^{\ast}$ and $f^{\ast}:=\min_{x\in{\mathbb{R}}^{d}}f(x)$. We denote $\Omega$ the underlying probability space over which any process and random variable of this paper is defined, and $\mathcal{P}(\Xi)$ some law over a metric space $\Xi$. For a random variable $X$ and any measurable function $\psi:\Xi\to{\mathbb{R}}$, we denote $\mathbb{E}_{\xi}[\psi(X,\xi)]:=\int_{\Xi}\psi(X,\xi)d\mathcal{P}(\xi)$, which is the conditional expectation of $\psi(X,\xi)$ with respect to $X$. $\mathcal{E}(1)$ denote the exponential distribution with parameter $1$, and $\Gamma(k,1)$ the Gamma distribution with shape parameter $k$ and rate $1$. For $a,b\in{\mathbb{R}}$, $a\wedge b=\min\{a,b\}$.
@@ 2.1 Geometrical Assumptions
It is immediate to check that under Assumption (PLμ), a critical point is also a global minimizer. It also ensures that $f$ has a quadratic growth, namely $f(x)-f^{\ast}\geq 0.5\mu\left\lVert x-x^{\ast}\right\rVert^{2}$, for all $x\in{\mathbb{R}}^{d}$ (bolte2017error; karimi2016linear). Gradient descent is known to make functions satisfying Assumptions (L-smooth) and (PLμ) converge linearly to their minimum, with a rate of order $\mathcal{O}((1-\mu/L)^{k}$) after $k$ iterations (bolte2017error; karimi2016linear). Importantly, momentum algorithms cannot significantly improve over this bound (PLlowerbound). In particular, the dependence on $\mu/L$ cannot be improved to $\sqrt{\mu/L}$ as in the strongly convex setting. To achieve such an acceleration, one can consider $(\tau,\mu)$-strong quasar convexity.
@@ 3 How to Use the Continuized Nesterov Framework
$x_{T_{k}}=x_{T_{k}^{-}}-\gamma\nabla f(x_{T_{k}^{-}},\xi_{k+1}),\quad z_{T_{k}}=z_{T_{k}^{-}}-\gamma^{\prime}\nabla f(z_{T_{k}^{-}},\xi_{k+1}).$
@@
$\left\{\begin{array}[]{ll}dx_{t}&=\eta(z_{t}-x_{t})dt-\gamma\int_{\Xi}\nabla f(x_{t},\xi)dN(t,\xi)\\ dz_{t}&=\eta^{\prime}(x_{t}-z_{t})dt-\gamma^{\prime}\int_{\Xi}\nabla f(x_{t},\xi)dN(t,\xi)\end{array}\right$. (17)
with the Poisson point measure $dN(t,\xi)=\sum_{k\geq 0}\delta_{T_{k},\xi_{k}}(dt,d\xi)$ on ${\mathbb{R}}_{+}\times\Xi$, which has intensity $dt\otimes\mathcal{P}$. By definition, the process is càdlàg, i.e. right-continuous with left limits. We assume $x_{0}=z_{0}$ is deterministic. Note that compared to the discussion of Section 1.2, we consider a constant set of parameters $(\eta,\eta^{\prime},\gamma,\gamma^{\prime})$.
@@ 3.1 Lyapunov Analysis with Stochastic Calculus
$dx_{t}=\zeta(x_{t})dt+\int_{\Xi}G(x_{t^{-}},\xi)dN(t,\xi)$
where $\zeta:{\mathbb{R}}^{d}\to{\mathbb{R}}^{d}$ is a Lipschitz function and $G:{\mathbb{R}}^{d}\times\Xi\to{\mathbb{R}}^{d}$ is measurable. Let $\varphi:{\mathbb{R}}^{d}\to{\mathbb{R}}$ be a $C^{1}$ function. Then,
@@ 3.2 Almost automatic discretization
Let $(x_{t},z_{t})_{t\geq 0}$ follow (17) with $\eta+\eta^{\prime}>0$. Define $\tilde{y}_{k}:=x_{T_{k+1}^{-}}$, $\tilde{x}_{k}:=x_{T_{k}}$ and $\tilde{z}_{k}:=z_{T_{k}}$ as evaluations of these processes. Then, $(\tilde{y}_{k},\tilde{x}_{k},\tilde{z}_{k})$ writes as a Nesterov algorithm in the form of (4) with stochastic parameters
@@
$\mathbb{E}\left[\liminf_{t\to+\infty}V_{t}\right]\leq\liminf_{t\to+\infty}\mathbb{E}\left[V_{t}\right]\leq\mathbb{E}\left[K_{0}\right].$ (23)
@@ 3.3 Application to Smooth, Polyak-Łojasiewicz Functions
Formalism setting   We start to set the formalism such that we can apply Proposition 3. Let $\overline{x}_{t}=(t,x_{t},z_{t})$ where $(x_{t},z_{t})_{t\geq 0}$ satisfy (17). It satisfies $d\overline{x}_{t}=\zeta(\overline{x}_{t})dt+G(\overline{x}_{t},\xi)dN(t,\xi)$, with
@@
Proof In Proposition 7, we proved (33), namely
@@
We deduce the result by applying Theorem 6 with $\varphi_{t}:=e^{\frac{\mu\gamma}{2}t}(f(x_{t})-f^{\ast})$, $K_{0}:=f(x_{0})-f^{\ast}$ and with the almost surely finite stopping time $T_{k}$. We conclude using Proposition 5, which ensures $\tilde{x}_{k}=x_{T_{k}}$.
@@
4 Improved Factor in Convergence Rate for Strongly Quasar Convex Functions and Better Trajectory-Wise Guarantees

### v2

@@ Abstract
In convex optimization, continuous-time counterparts have been a fruitful tool for analyzing momentum algorithms. Fewer such examples are available when the function to minimize is non-convex. In several cases, discrepancies arise between the existing discrete-time results — namely those obtained for momentum algorithms — and their continuous-time counterparts, with the latter typically yielding stronger guarantees. We argue that the continuized framework (even2021continuized), mixing continuous and discrete components, can tighten the gap between known continuous and discrete results. This framework relies on computations akin to standard Lyapunov analyses, from which are deduced convergence bounds for an algorithm that can be written as a Nesterov momentum algorithm with stochastic parameters. In this work, we extend the range of applicability of the continuized framework, e.g. by allowing it to handle non-smooth Lyapunov functions. We then strengthen its trajectory-wise guarantees for linear convergence rate, deriving finite time bounds with high probability and asymptotic almost sure bounds. We apply this framework to the non-convex class of strongly quasar-convex functions. Adapting continuous-time results that have weaker discrete equivalents to the continuized method, we improve by a constant factor the known convergence rate, and relax the existing assumptions on the set of minimizers.
@@ 1.1 ODEs Momentum and their Limit for Algorithm Design
Note that $\{\tilde{x}_{k}\}_{k\in{\mathbb{N}}}$ reduces to gradient descent for the choice $\alpha_{k}=1$, $\forall k\in{\mathbb{N}}$. As previously mentioned, it is known to accelerate over gradient descent in many settings in convex optimization. In particular, for $\mu$-strongly convex functions with $L$-Lipschitz gradient, $\{\tilde{x}_{k}\}_{k\in{\mathbb{N}}}$ generated by (4) properly parameterized ensures $f(\tilde{x}_{k})-f^{\ast}=\mathcal{O}((1-\sqrt{\mu/L})^{k})$ (nesterovbook). This should be compared with the bound $\mathcal{O}((1-\mu/L)^{k})$ using gradient descent, which can be significantly worse as typically $\mu\ll L$ (this is particularly true in the high dimensional setting). It is thus common to mention this gain of a squared dependence on $\mu/L$ as acceleration.
@@
Both rates correspond to linear convergence, with rates $\sqrt{\mu}$ and $\sqrt{\mu/L}$, respectively. In this sense, the continuous-time result is transferred to the discrete setting. The $L$-Lipschitzness of the gradient is a key property to enable this transfer. In particular, the accelerated rate $(1-\sqrt{\mu/L})^{k}$ can be achieved if we take $s$ in (4) large enough, namely $s=1/L$. This has been extensively used in the convex setting where similar convergence results are obtained for $(x_{t})_{t\geq 0}$ and $\{x_{n}\}_{n\geq 0}$, see e.g. (suboydcandes; siegel2021accelerated; attouch2020firstorder; shi2022understanding; aujdossrondPL; hyppo; li2024linear). However, this approach may suffer from two drawbacks.
(i) Convergence properties do not transfer   The convergence properties of $(x_{t})_{t\geq 0}$ may not transfer to $\{x_{n}\}_{n\geq 0}$ when $f$ is not convex. In the case of $L$-smooth functions, if the total energy decreases when applied to the solution of the continuous Heavy Ball equation, it is not the case for the discrete counterpart, see (jin2018accelerated, Section A.1). For functions with Lipschitz gradient and Hessian, a convergence result exists for the continuous Heavy Ball equation, but needs additional restrictions to deduce a discrete counterpart, see the first version of (okamura2024primitive). In the case of strongly quasar-convex functions, in order to select a stepsize large enough to achieve the best known convergence rate with $\{\tilde{x}_{k}\}_{k\in{\mathbb{N}}}$ generated by (4), the non-convexity has to be restricted (hermant2024study; guptanesterov). Such restrictions are not necessary when considering $(x_{t})_{t\geq 0}$ generated by (7). Using side by side comparisons of discrete and continuous Lyapunov functions, Section 4.2 in (hermant2024study) precises how non-convexity hurts the discretization process in this case.
@@ 1.2 A Promising Alternative: Continuized Nesterov
At $t=T_{i}$, the process jumps by performing a gradient step
@@
$\left\{\begin{array}[]{ll}dx_{t}&=\eta_{t}(z_{t}-x_{t})dt-\gamma_{t^{-}}\nabla f(x_{t^{-}})N(dt)\\ dz_{t}&=\eta^{\prime}_{t}(x_{t}-z_{t})dt-\gamma^{\prime}_{t^{-}}\nabla f(x_{t^{-}})N(dt)\end{array}\right$. (14)
where $N(dt)=\sum_{k\geq 0}\delta_{T_{k}}(dt)$ is a Poisson point measure. Intuitively, since $\mathbb{E}\left[dN_{t}\right]=dt$, the mean process $(\mathbb{E}\left[x_{t}\right],\mathbb{E}\left[z_{t}\right])_{t\geq 0}$ is smooth, see Figure 1. In particular, if $\nabla f$ is linear, it can be written exactly in the form of (7). Consequently, individual realizations of the jump-driven process (14) provide a stochastic approximation of a much smoother expected trajectory. It implies that, for a given Lyapunov function $E_{t}:=\varphi(x_{t},z_{t},t)$, the analysis of $\mathbb{E}\left[E_{t}\right]$ is close to that arising in the standard continuous setting, such as in the study of (7).
@@
A non-convex application: strong quasar-convexity   Strong quasar-convexity is a non-convex relaxation of strong convexity, that has received increasing attention in recent years (hinder2020near; gower2021sgd; wang2023continuizedaccelerationquasarconvex; fu2023accelerated; alimisis2024characterization; hermant2024study; chenefficient2025; guptanesterov; lara2025delayed; de2025extending; farzin2025minimisation). Compared to gradient descent, better convergence bounds exist for Nesterov momentum when minimizing these functions. The first of these accelerated results mixes the momentum mechanism with a subroutine that solves an optimization sub-problem (quasarconvexold1; quasarconvexold2), reduced to a line-search procedure which adds a logarithmic factor in the complexity in (hinder2020near). Avoiding such a subroutine requires to restrict the class of function, excluding functions that are too non-convex (hermant2024study; guptanesterov). The continuized method has been applied to strongly quasar-convex functions (wang2023continuizedaccelerationquasarconvex), recovering the convergence rate of the classical Nesterov momentum algorithm. Strikingly, it does so without requiring a line-search procedure or restricting the level of non-convexity. This indicates that the method is capable of achieving results that classical momentum algorithms have not been able to obtain so far, and we show in this work that it can go further.
@@ 1.3 Limits of the Framework and Contributions
(iii) Adapting results coming from the classic ODE framework literature, we improve the best known convergence rate by a constant factor for smooth strongly quasar-convex functions. We further relax the existing assumption on the set of minimizers for this class of functions. Our results hold when considering unbiased stochastic estimates of the gradients, as long as a strong growth condition is verified.
@@ 2 Assumptions and Notations
We denote $\mathcal{X}^{\ast}:=\arg\min_{{\mathbb{R}}^{d}}f(x)$, $x^{\ast}$ any element of $\mathcal{X}^{\ast}$ and $f^{\ast}:=\min_{x\in{\mathbb{R}}^{d}}f(x)$. We denote $\Omega$ the underlying probability space over which any process and random variable of this paper is defined, and $\mathcal{P}(\Xi)$ some law over a metric space $\Xi$. For a random variable $X$ and any measurable function $\psi:\Xi\to{\mathbb{R}}$, we denote $\mathbb{E}_{\xi}[\psi(X,\xi)]:=\int_{\Xi}\psi(X,\xi)\mathcal{P}(d\xi)$, which is the conditional expectation of $\psi(X,\xi)$ with respect to $X$. $\mathcal{E}(1)$ denotes the exponential distribution with parameter $1$, and $\Gamma(k,1)$ the Gamma distribution with shape parameter $k$ and rate $1$. For $a,b\in{\mathbb{R}}$, $a\wedge b=\min\{a,b\}$.
@@ 2.1 Geometrical Assumptions
It is immediate to check that under Assumption (PLμ), a critical point is also a global minimizer. It also ensures that $f$ has a quadratic growth, namely $f(x)-f^{\ast}\geq 0.5\mu\left\lVert x-x^{\ast}\right\rVert^{2}$, for all $x\in{\mathbb{R}}^{d}$ (bolte2017error; karimi2016linear). Gradient descent is known to make functions satisfying Assumptions (L-smooth) and (PLμ) converge linearly to their minimum, with a rate of order $\mathcal{O}((1-\mu/L)^{k}$) after $k$ iterations (bolte2017error; karimi2016linear). Importantly, momentum algorithms cannot significantly improve over this bound (PLlowerbound). In particular, the dependence on $\mu/L$ cannot be improved to $\sqrt{\mu/L}$ as in the strongly convex setting. To achieve such an acceleration, one can consider $(\tau,\mu)$-strong quasar-convexity.
@@ 3 How to Use the Continuized Nesterov Framework
$x_{T_{k}}=x_{T_{k}^{-}}-\gamma\nabla f(x_{T_{k}^{-}},\xi_{k}),\quad z_{T_{k}}=z_{T_{k}^{-}}-\gamma^{\prime}\nabla f(z_{T_{k}^{-}},\xi_{k}).$
@@
$\left\{\begin{array}[]{ll}dx_{t}&=\eta(z_{t}-x_{t})dt-\gamma\int_{\Xi}\nabla f(x_{t^{-}},\xi)N(dt,d\xi)\\ dz_{t}&=\eta^{\prime}(x_{t}-z_{t})dt-\gamma^{\prime}\int_{\Xi}\nabla f(x_{t^{-}},\xi)N(dt,d\xi)\end{array}\right$. (17)
with the Poisson point measure $N(dt,d\xi)=\sum_{k\geq 0}\delta_{(T_{k},\xi_{k})}(dt,d\xi)$ on ${\mathbb{R}}_{+}\times\Xi$, which has intensity $dt\otimes\mathcal{P}$. By definition, the process is càdlàg, i.e. right-continuous with left limits. We assume $x_{0}=z_{0}$ is deterministic. Note that compared to the discussion of Section 1.2, we consider a constant set of parameters $(\eta,\eta^{\prime},\gamma,\gamma^{\prime})$.
@@ 3.1 Lyapunov Analysis with Stochastic Calculus
$dx_{t}=\zeta(x_{t})dt+\int_{\Xi}G(x_{t^{-}},\xi)N(dt,d\xi)$
where $\zeta:{\mathbb{R}}^{d}\to{\mathbb{R}}^{d}$ is a locally Lipschitz function and $G:{\mathbb{R}}^{d}\times\Xi\to{\mathbb{R}}^{d}$ is measurable. Let $\varphi:{\mathbb{R}}^{d}\to{\mathbb{R}}$ be with Lipschitz gradient. Then,
@@ 3.2 Almost Automatic Discretization
Let $(x_{t},z_{t})_{t\geq 0}$ follow (17) with $\eta+\eta^{\prime}>0$ and with underlying jumping times $\{T_{k}\}_{k\in{\mathbb{N}}^{\ast}}$. Define $\tilde{y}_{k}:=x_{T_{k+1}^{-}}$, $\tilde{x}_{k+1}:=x_{T_{k+1}}$ and $\tilde{z}_{k+1}:=z_{T_{k+1}}$ as evaluations of this process. Then, $(\tilde{y}_{k},\tilde{x}_{k},\tilde{z}_{k})$ writes as a Nesterov algorithm in the form of (4) with stochastic parameters
@@
$\mathbb{E}\left[\liminf_{t\to+\infty}\>V_{t}\right]\leq\liminf_{t\to+\infty}\>\mathbb{E}\left[V_{t}\right]\leq\mathbb{E}\left[K_{0}\right].$ (23)
@@ 3.3 Application to Smooth, Polyak-Łojasiewicz Functions
Formalism Setting   We start to set the formalism such that we can apply Proposition 3. Let $\overline{x}_{t}=(t,x_{t},z_{t})$ where $(x_{t},z_{t})_{t\geq 0}$ satisfy (17). It satisfies $d\overline{x}_{t}=\zeta(\overline{x}_{t})dt+G(\overline{x}_{t},\xi)N(dt,d\xi)$, with
@@
Proof To prove Proposition 7, we established that for $(x_{t},z_{t})_{t\geq 0}$ solution of (17) with $\gamma\leq\frac{1}{\rho L}$, $\gamma^{\prime}=\gamma+\sqrt{\frac{\gamma}{2\rho}}$, $\eta=\sqrt{\frac{\gamma}{2\rho}}$ and $\eta^{\prime}=\frac{\mu\gamma}{2}-\sqrt{\frac{\gamma}{2\rho}}$, we have (33), namely
@@
Applying Theorem 6 with $\varphi_{t}:=e^{\frac{\mu\gamma}{2}t}(f(x_{t})-f^{\ast})$, $K_{0}:=f(x_{0})-f^{\ast}$ and with the almost surely finite stopping time $T_{k}$ yields
We conclude using Proposition 5, which states that if $(x_{t},z_{t})_{t\geq 0}$ is generated with the stopping times $\{T_{k}\}_{k\in{\mathbb{N}}^{\ast}}$, then $\tilde{x}_{k}=x_{T_{k}}$, for $\{\tilde{x}_{k}\}_{k\in{\mathbb{N}}}$ generated by (22) with choice of parameters as stated.
@@
4 Improved Factor in Convergence Rate for Strongly quasar-convex Functions and Better Trajectory-Wise Guarantees